# KP-PINNs: Kernel Packet Accelerated Physics Informed Neural Networks

### v1

@@ Abstract
Differential equations are involved in modeling many engineering problems. Many efforts have been devoted to solving differential equations. Due to the flexibility of neural networks, Physics Informed Neural Networks (PINNs) have recently been proposed to solve complex differential equations and have demonstrated superior performance in many applications. While the L2 loss function is usually a default choice in PINNs, it has been shown that the corresponding numerical solution is incorrect and unstable for some complex equations. In this work, we propose a new PINNs framework named Kernel Packet accelerated PINNs (KP-PINNs), which gives a new expression of the loss function using the reproducing kernel Hilbert space (RKHS) norm and uses the Kernel Packet (KP) method to accelerate the computation. Theoretical results show that KP-PINNs can be stable across various differential equations. Numerical experiments illustrate that KP-PINNs can solve differential equations effectively and efficiently. This framework provides a promising direction for improving the stability and accuracy of PINNs-based solvers in scientific computing.
@@ 3 KP Accelerated PINNs
By kernel packet method introduced in Section 2.3, ${\rm Loss}_{\mathcal{N}_{\Phi_{T}^{\nu}}}$ can be easily computed.Some well-established optimization algorithms, such as Adam or L-BFGS, can be used to minimize the loss effectively. The KP-PINNs algorithm is summarized in Algorithm 1. For the inverse problem, comparing (7) and (8), it can be seen that the loss function of the inverse problem has more ${\rm Loss}_{\mathcal{D}}(u)$ part. For this part, the computation of the RKHS norm is similar to the analysis above. In this way, the KP-PINNs algorithm can effectively solve both forward and inverse problems.
@@ 5.1 Stiff Equation
The Stiff equation Wanner and Hairer (1996) exhibits steep gradients and is typically handled via time-scale decomposition to reduce reaction-related errors. It is defined by
@@ 5.2 Helmholtz Equation
Helmholtz equation is an elliptic PDE in electromagnetism that describes the electromagnetic wave propagation with
@@ 6 Conclusion
This work proposes KP-PINNs, an efficient method for solving differential equations, with stability guarantees for second-order linear elliptic and HJB equations under certain conditions. Numerical experiments demonstrate that KP-PINNs outperform $L_{2}$-PINNs and Sobolev-PINNs for both forward and inverse problems, significantly accelerating the computation of the kernel matrix inverse $\mathbf{K}$ through the KP technique. Future directions include extending KP-PINNs to handle sparse or irregular data, optimizing neural network architecture, and refining kernel selection in the loss function to improve accuracy and robustness. These advancements aim to enhance KP-PINNs’ versatility in solving complex differential equations.

### v2

@@ Abstract
Differential equations are involved in modeling many engineering problems. Many efforts have been devoted to solving differential equations. Due to the flexibility of neural networks, Physics Informed Neural Networks (PINNs) have recently been proposed to solve complex differential equations and have demonstrated superior performance in many applications. While the $L_{2}$ loss function is usually a default choice in PINNs, it has been shown that the corresponding numerical solution is incorrect and unstable for some complex equations. In this work, we propose a new PINNs framework named Kernel Packet accelerated PINNs (KP-PINNs), which gives a new expression of the loss function using the Reproducing Kernel Hilbert Space (RKHS) norm and uses the Kernel Packet (KP) method to accelerate the computation. Theoretical results show that KP-PINNs can be stable across various differential equations. Numerical experiments illustrate that KP-PINNs can solve differential equations effectively and efficiently. This framework provides a promising direction for improving the stability and accuracy of PINNs-based solvers in scientific computing.
@@ 3 KP Accelerated PINNs
By KP method introduced in Section 2.3, ${\rm Loss}_{\mathcal{N}_{\Phi_{T}^{\nu}}}$ can be easily computed.Some well-established optimization algorithms, such as Adam or L-BFGS, can be used to minimize the loss effectively. The KP-PINNs algorithm is summarized in Algorithm 1. For the inverse problem, comparing (7) and (8), it can be seen that the loss function of the inverse problem has more ${\rm Loss}_{\mathcal{D}}(u)$ part. For this part, the computation of the RKHS norm is similar to the analysis above. In this way, the KP-PINNs algorithm can effectively solve both forward and inverse problems.
@@ 5.1 Stiff Equation
The Stiff equation Wanner and Hairer (1996) exhibits steep gradients caused by fast dynamics, and is commonly treated with time-scale decomposition to reduce numerical and reaction-related errors. It is defined by
@@ 5.2 Helmholtz Equation
The Helmholtz equation is a second-order elliptic PDE commonly encountered in electromagnetism that describes the spatial behavior of electromagnetic wave propagation. It is defined by
@@ 6 Conclusion
This work proposes KP-PINNs, an efficient method for solving differential equations, with stability guarantees for second-order linear elliptic and HJB equations under certain conditions. Numerical experiments demonstrate that KP-PINNs outperform $L_{2}$-PINNs and Sobolev-PINNs for both forward and inverse problems, significantly accelerating the computation of the inverse of kernel matrix $\mathbf{K}$ through the KP technique. Future directions include extending KP-PINNs to handle sparse or irregular data, optimizing neural network architecture, and refining kernel selection in the loss function to improve accuracy and robustness. These advancements aim to enhance KP-PINNs’ versatility in solving complex differential equations.